OpenStudy (anonymous):

find all solutions of x = 3/x + 1/2

OpenStudy (anonymous):

The degree of x is only 1, so there is just 1 solution. x = 3/x + 1/2 x-3/x = 1/2 (x^2-3)/x = 1/2 x^2-3=1/2x x^2-1/2x-3=0 Now factor the solutions :)